# MetaSel: A Test Selection Approach for Fine-tuned DNN Models

### v2

@@ VI-C RQ3: MetaSel’s effectiveness across varying levels of distribution shifts
To address the first two research questions, we conducted our experiments on subjects representing a medium level of distribution shift (severity level = 3), as mentioned earlier. To answer this research question, we further extend our analysis by repeating the experiments from RQ1 using two additional severity levels representing both weaker (severity level = 2) and stronger (severity level = 4) distribution shifts. Such analysis offers valuable insights into MetaSel’s robustness to real-world scenarios where distribution shifts can vary in severity. To this end, we utilized the same models and source inputs detailed in Table II, and employed the same corruption types we used in our earlier experiments. Similar to RQ1, we calculated the TRC achieved by MetaSel and each baseline using Equation (7) across the same range of budget sizes. As detailed below, the results demonstrate that MetaSel consistently and significantly outperforms all baselines, achieving higher TRC values across all severity levels, subjects, and selection budgets. To assess the statistical significance of these performance improvements, we once again conducted Wilcoxon signed-rank tests with a significance level of $\alpha=0.05$. Results show that all p-values are below 0.05, confirming MetaSel’s consistently superior performance across varying severity levels of distribution shift. To provide a comprehensive report, we have included the full results of our statistical analysis, detailed TRC values, and TRC curves for MetaSel and all baseline approaches across all severity levels of distribution shift, subjects, and selection budgets in our replication package [22].
@@
Consistent with our earlier findings on subjects with medium levels of distribution shift, the second-best performing approach after MetaSel also varies across subjects, selection budgets, and levels of distribution shift. These findings further confirm that MetaSel is reliably outperforming baselines, regardless of subjects, selection budgets, and distribution shift levels. The three most frequently occurring second-best baselines, in subjects with both weaker and stronger severity levels of distribution shifts are once again Meta-model, DATIS, and NNS, similar to what was observed in RQ1. Consequently, we only report in Table V the average improvement percentages provided by MetaSel over these three baselines. Similar to RQ1, we focused on highly constrained budget sizes of 1%, 3%, 5%, and 10% of the target test set’s size. Consistent with our earlier findings in RQ1 on subjects with a medium level of distribution shift, MetaSel yields practically significant improvements relative to the best-performing baselines in subjects with both weaker and stronger shift severity levels. For subjects with weaker distribution shifts (severity level = 2), MetaSel achieves average improvement percentages ranging from 30.97% to 45.07% over Meta-model, 36.74% to 39.17% over DATIS, and 38.87% to 56.18% over NNS. Similarly, for subjects with stronger distribution shifts (severity level = 4), MetaSel maintains its superiority, achieving average improvement percentages of 28.46% to 41.23% over Meta-model, 31.55% to 38.42% over DATIS, and 33.42% to 53.46% over NNS. These results highlight that MetaSel consistently provides a practically significant improvement in misclassification detection across a large distribution shift severity range. They thus confirm MetaSel’s wide applicability across diverse distribution shift situations in practice.
@@
VI-D Threats to Validity
@@ VII Related Work
Test selection for DNNs. In recent years, several test selection approaches have been proposed for DNN models aiming to identify unlabeled inputs with a higher probability of being mispredicted by the model. These approaches are broadly classified into black-box and white-box approaches, depending on whether they require access to the internals of the DNN model [2]. Neuron coverage (NC) metrics [38, 39], a class of the white-box approaches that have been proposed for DNNs, are inspired by the success of code coverage in traditional software. However, studies have consistently shown that these criteria lack a significant correlation with the number of mispredicted inputs in a test set [40, 41, 42].
Another early contribution in this category are SA metrics proposed by Kim et al. [3]. These metrics measure how surprising a test input is with respect to the DNN model’s training input set. Test inputs are then ranked in ascending order based on their SA scores, with more surprising inputs given higher priority for testing. Our results confirm earlier findings [6, 9] that DSA delivers better performance compared to the other alternative SA metrics. However, MetaSel consistently outperforms all SA metrics across all subjects and distribution shift levels.
Black-box approaches rely on the output of MUT to guide test selection. For instance, ATS introduced by Gao et al. [13], is a diversity-driven approach that leverages the variations in model outputs as a behavioral diversity metric. By focusing on these differences, ATS aims to identify and cover a broader range of potential failures. However, ATS involves computing distances between inputs and has been demonstrated to be one of the most computationally intensive approaches [14, 8], making it impractical particularly when dealing with large input sets. One of the most effective and widely used black-box approaches for test input prioritization is probability-based uncertainty metrics [5, 6, 1, 7]. Their simplicity and reliance solely on MUT’s output probability vector make them highly efficient. In training MetaSel, however, we utilize the raw outputs of the logit layer instead of the probability vector, as logits enable more fine-grained comparisons between the pre-trained and fine-tuned DNN models. Our preliminary study on feature selection further validates this choice, as detailed in Section IV-A. Our results confirm the efficiency and effectiveness of Probability-based uncertainty metrics. They also demonstrate that MetaSel consistently outperforms these approaches across subjects.
@@
Our results demonstrate that MetaSel consistently outperforms all the investigated SOTA general test selection approaches including white-box, black-box, and learning-based baselines across all subjects. This highlights the significant potential of leveraging additional contextual information, when available, to improve test input prioritization and selection. Furthermore, our empirical study on 10 baselines reveals that the most frequent second-best approach after MetaSel is Meta-model, a learning-based approach. This further highlights the growing importance and effectiveness of learning-based techniques, suggesting that these approaches can play a pivotal role in enhancing test input selection effectiveness.
@@ VIII Conclusion
In this paper, we introduce MetaSel, a test input prioritization and selection method specifically designed for fine-tuned DNN models. Our approach leverages information about the relationship between the fine-tuned model and its pre-trained counterpart to estimate misclassification probabilities for unlabeled test inputs. MetaSel relies on the logits of both the pre-trained and fine-tuned models, their differences, and the comparison of their predicted output labels. Additionally, MetaSel employs a measure of how well an input aligns with the data distributions on which both the pre-trained and fine-tuned models were trained. Our empirical analysis, comparing MetaSel against 10 SOTA test selection approaches, demonstrates that MetaSel consistently outperforms all baselines, particularly under tightly constrained labeling budgets, which is especially important when testing fine-tuned DNN models. The results show that MetaSel is capable of detecting a significantly higher number of misclassified inputs within the same selection budget. In our extensive empirical study, involving 68 subjects across varying levels of distribution shift severity between pre-trained and fine-tuned models, MetaSel’s performance remained consistent regardless of subject, selection budget, and distribution shift severity level. Further, the second-best performing approach after MetaSel varies across different subjects and selection budgets, further confirming that MetaSel is the best solution. Additionally, MetaSel not only achieves higher TRC median and average but also maintains significantly lower variability in TRC values than alternative baselines across all subjects. Last, when assessing the computational efficiency of MetaSel in comparison to baselines, we demonstrate MetaSel’s practicality and cost-effectiveness, even when applied to large input sets and complex DNN models. Although MetaSel is inherently adaptable and applicable to a wide range of input types, we plan to extend our investigation to include input sets from diverse application domains to determine how well MetaSel maintains its effectiveness across different data domains.

### v4

@@ VI-C RQ3: MetaSel’s effectiveness across varying levels of distribution shifts
To address the first two research questions, we conducted our experiments on subjects representing a medium level of distribution shift (severity level = 3), as mentioned earlier. To answer this research question, we further extend our analysis by repeating the experiments from RQ1 using two additional severity levels representing both weaker (severity level = 2) and stronger (severity level = 4) distribution shifts. Such analysis offers valuable insights into MetaSel’s robustness to real-world scenarios where distribution shifts can vary in severity. To this end, we utilized the same models and source inputs detailed in Table II, and employed the same corruption types we used in our earlier experiments.
Similar to RQ1, we calculated the TRC achieved by MetaSel and each baseline using Equation (7) across the same range of budget sizes. As detailed below, the results demonstrate that MetaSel consistently and significantly outperforms all baselines, achieving higher TRC values across all severity levels, subjects, and selection budgets. To assess the statistical significance of these performance improvements, we once again conducted Wilcoxon signed-rank tests with a significance level of $\alpha=0.05$. Results show that all p-values are below 0.05, confirming MetaSel’s consistently superior performance across varying severity levels of distribution shift. To provide a comprehensive report, we have included the full results of our statistical analysis, detailed TRC values, and TRC curves for MetaSel and all baseline approaches across all severity levels of distribution shift, subjects, and selection budgets in our replication package [22].
@@
Consistent with our earlier findings on subjects with medium levels of distribution shift, the second-best performing approach after MetaSel also varies across subjects, selection budgets, and levels of distribution shift. These findings further confirm that MetaSel is reliably outperforming baselines, regardless of subjects, selection budgets, and distribution shift levels. The three most frequently occurring second-best baselines, in subjects with both weaker and stronger severity levels of distribution shifts, are once again Meta-model, DATIS, and NNS, similar to what was observed in RQ1. Consequently, we only report in Table V the average improvement percentages provided by MetaSel over these three baselines. Similar to RQ1, we focused on highly constrained budget sizes of 1%, 3%, 5%, and 10% of the target test set’s size. Consistent with our earlier findings in RQ1 on subjects with a medium level of distribution shift, MetaSel yields practically significant improvements relative to the best-performing baselines in subjects with both weaker and stronger shift severity levels. For subjects with weaker distribution shifts (severity level = 2), MetaSel achieves average improvement percentages ranging from 30.97% to 45.07% over Meta-model, 36.74% to 39.17% over DATIS, and 38.87% to 56.18% over NNS. Similarly, for subjects with stronger distribution shifts (severity level = 4), MetaSel maintains its superiority, achieving average improvement percentages of 28.46% to 41.23% over Meta-model, 31.55% to 38.42% over DATIS, and 33.42% to 53.46% over NNS. These results highlight that MetaSel consistently provides a practically significant improvement in misclassification detection across a large distribution shift severity range. They thus confirm MetaSel’s wide applicability across diverse distribution shift situations in practice.
@@
VI-D RQ4: Contribution of MetaSel’s training features
To address this research question, we conducted a comprehensive ablation study comparing MetaSel with seven ablated variants created by both single-feature and grouped-feature ablation strategies.
As described in Section IV-A, MetaSel’s training feature set comprises six features obtained from both source and target models, including the logit vector of the source model ($L_{S}$), the logit vector of the target model ($L_{T}$), the absolute difference between the two logits ($L_{S}-L_{T}$), the ODIN score based on $M_{S}$ ($ODIN_{S}$), the ODIN score based on $M_{T}$ ($ODIN_{T}$), and a binary indicator of differential testing outcome (1 if $M_{S}$ and $M_{T}$ predict the same class, 0 otherwise). We systematically constructed seven variants of MetaSel by excluding one or more features from this set.
The variants are as follows:
  • V1: excludes the difference between logits ($L_{S}-L_{T}$),
  • V2: excludes the differential testing outcome,
  • V3: excludes $ODIN_{S}$,
  • V4: excludes both ODIN scores ($ODIN_{S}$ and $ODIN_{T}$),
  • V5: excludes $L_{S}$ and thus $L_{S}-L_{T}$,
  • V6: excludes logits ($L_{S}$, $L_{T}$, and thus $L_{S}-L_{T}$),
  • V7: excludes all information from $M_{S}$ and thus includes only $L_{T}$ and $ODIN_{T}$.
Variants such as V1, V2, and V3 represent single-feature ablation, where only one feature is removed to isolate its contribution. In contrast, variants such as V4, V5, V6, and V7 represent grouped feature ablation, where multiple features are removed together based on their origin or semantic grouping (e.g., features originated from the source model or OOD-based features). By comparing the performance of these variants to the original MetaSel model, we evaluate the contribution of each feature or group of features to MetaSel’s ability to effectively select misclassified target test inputs.
Note that we performed our experiments only on subjects representing a medium level of distribution shift (i.e., severity level 3). This choice is based on our earlier findings from RQ3, which demonstrated that MetaSel consistently maintains its effectiveness across different levels of distribution shift severity. To answer this question, we compare the TRC results achieved by MetaSel with those of variants across all subjects (i.e., input sets, DNN models, and corruptions presented in Table V). We compare the TRC distributions achieved by MetaSel and all variants under highly constrained budgets of 1%, 3%, 5%, and 10% of the target test set, illustrated in Figure 6. To provide a comprehensive insight, we report the complete TRC curves achieved by all variants in our replication package [22].
$b$ = 1%
$b$ = 3%
$b$ = 5%
$b$ = 10%
Figure 6: TRC distribution achieved by MetaSel and its seven variants across all subjects under highly constrained selection budgets, across subjects representing a medium severity level of distribution shift (severity level = 3)
Each box plot in Figure 6, presents the distribution of TRC achieved by MetaSel and its ablated variants under a given selection budget $b$. MetaSel consistently achieves the highest TRC median and average values across all budget sizes, validating its superior effectiveness in identifying misclassified inputs compared to any of its variants. MetaSel’s lower TRC variance compared to its variants further supports its reliability across diverse subjects. These results confirm that removing any of the training features leads to significant performance degradation, thereby validating the essential contribution of each individual feature used in training MetaSel.
Answer to RQ4: Results of our comprehensive ablation study comparing MetaSel with seven ablated variants across all subjects confirm that all features selected for constructing MetaSel significantly contribute to enhancing its performance.
VI-E Threats to Validity
@@ VII Related Work
Test selection for DNNs. In recent years, several test selection approaches have been proposed for DNN models aiming to identify unlabeled inputs with a higher probability of being mispredicted by the model. These approaches are broadly classified into black-box and white-box approaches, depending on whether they require access to the internals of the DNN model [3]. Neuron coverage (NC) metrics [42, 43], a class of white-box approaches that have been proposed for DNNs, are inspired by the success of code coverage in traditional software. However, studies have consistently shown that these criteria lack a significant correlation with the number of mispredicted inputs in a test set [44, 45, 46].
Another early contribution in this category is SA metrics proposed by Kim et al. [4]. These metrics measure how surprising a test input is with respect to the DNN model’s training input set. Test inputs are then ranked in ascending order based on their SA scores, with more surprising inputs given higher priority for testing. Our results confirm earlier findings [7, 10] that DSA delivers better performance compared to the other alternative SA metrics. However, MetaSel consistently outperforms all SA metrics across all subjects and distribution shift levels.
Black-box approaches rely on the output of MUT to guide test selection. For instance, ATS introduced by Gao et al. [17], is a diversity-driven approach that leverages the variations in model outputs as a behavioral diversity metric. By focusing on these differences, ATS aims to identify and cover a broader range of potential failures. However, ATS involves computing distances between inputs and has been demonstrated to be one of the most computationally intensive approaches [15, 9], making it impractical, particularly when dealing with large input sets. One of the most effective and widely used black-box approaches for test input prioritization is probability-based uncertainty metrics [6, 7, 2, 8]. Their simplicity and reliance solely on MUT’s output probability vector make them highly efficient. In training MetaSel, however, we utilize the raw outputs of the logit layer instead of the probability vector, as logits enable more fine-grained comparisons between the pre-trained and fine-tuned DNN models. Our comprehensive ablation study on selected features further validates this choice, as detailed in Section VI-D. Our results confirm the efficiency and effectiveness of Probability-based uncertainty metrics. They also demonstrate that MetaSel consistently outperforms these approaches across subjects.
@@
RTS proposed by Sun et al. [20] combines probability-based uncertainty with diversity for test prioritization. Like other diversity-based methods, it relies on pairwise similarity computations (based on SSIM similarity [21]), resulting in high computational overhead. As shown in our results, RTS is the most time-consuming approach, taking approximately 4 to 15 times longer than MetaSel to prioritize the entire test set across our subjects. Moreover, RTS is consistently outperformed by MetaSel in terms of effectiveness and never appears among the second-best baselines.
@@
Our results demonstrate that MetaSel consistently outperforms all the investigated SOTA general test selection approaches including white-box, black-box, and learning-based baselines across all subjects. This highlights the significant potential of leveraging additional contextual information, when available, to improve test input prioritization and selection. Furthermore, our empirical study on 11 baselines reveals that the most frequent second-best approach after MetaSel is Meta-model, a learning-based approach. This further highlights the growing importance and effectiveness of learning-based techniques, suggesting that these approaches can play a pivotal role in enhancing test input selection effectiveness.
@@ VIII Conclusion
In this paper, we introduce MetaSel, a test input prioritization and selection method specifically designed for fine-tuned DNN models. Our approach leverages information about the relationship between the fine-tuned model and its pre-trained counterpart to estimate misclassification probabilities for unlabeled test inputs. MetaSel relies on the logits of both the pre-trained and fine-tuned models, their differences, and the comparison of their predicted output labels. Additionally, MetaSel employs a measure of how well an input aligns with the data distributions on which both the pre-trained and fine-tuned models were trained. Our empirical analysis, comparing MetaSel against 11 SOTA test selection approaches, demonstrates that MetaSel consistently outperforms all baselines, particularly under tightly constrained labeling budgets, which is especially important when testing fine-tuned DNN models. The results show that MetaSel is capable of detecting a significantly higher number of misclassified inputs within the same selection budget. In our extensive empirical study, involving 68 subjects across varying levels of distribution shift severity between pre-trained and fine-tuned models, MetaSel’s performance remained consistent regardless of subject, selection budget, and distribution shift severity level. Further, the second-best performing approach after MetaSel varies across different subjects and selection budgets, further confirming that MetaSel is the best solution. Additionally, MetaSel not only achieves higher TRC median and average but also maintains significantly lower variability in TRC values than alternative baselines across all subjects. Last, when assessing the computational efficiency of MetaSel in comparison to baselines, we demonstrate MetaSel’s practicality and cost-effectiveness, even when applied to large input sets and complex DNN models. Although MetaSel is inherently adaptable and applicable to a wide range of input types, we plan to extend our investigation to include input sets from diverse application domains, such as text inputs, to determine how well MetaSel maintains its effectiveness across different data domains. Another direction for our future work is to explore the potential of extending and adapting MetaSel to transfer learning scenarios beyond covariate shift, encompassing different but related source and target domains that still result in successful fine-tuning. Although MetaSel in its original design may not be directly applicable to all such scenarios, adaptation strategies could enable its broader applicability. For instance, when the source and target input sets feature different output classes, a mapping between these classes would enable the extraction of MetaSel’s features, e.g., calculating the differential testing outcome and measuring the differences between the source and target logit vectors.